OpenStudy (anonymous):

Solve for x. 1.2^(x)=(0.5)^x If you could show the steps that would help!!:)

OpenStudy (amistre64):

logs help

OpenStudy (amistre64):

a^x = b^x x loga = x logb x loga - x logb = 0 x (loga - logb) = 0

OpenStudy (amistre64):

either a/b = 1, or x is equla to ?

OpenStudy (anonymous):

Thank you!